# Minimal thermodynamic cost of computing with circuits

### v1

@@ Abstract
All digital devices have components that implement Boolean functions, mapping that component’s input to its output. However, any fixed Boolean function can be implemented by an infinite number of circuits, all of which vary in their resource costs. This has given rise to the field of circuit complexity theory, which studies the minimal resource cost to implement a given Boolean function with any circuit. Traditionally, circuit complexity theory has focused on the resource costs of a circuit’s size (its number of gates) and depth (the longest path length from the circuit’s input to its output). In this paper, we extend circuit complexity theory by investigating the minimal thermodynamic cost of a circuit’s operation. We do this by using the mismatch cost of a given circuit that is run multiple times in a row to calculate a lower bound on the entropy production (EP) incurred in each such run of the circuit. Specifically, we derive conditions for mismatch cost to be proportional to the size of a circuit, and conditions for them to diverge. We also use our results to compare the thermodynamic costs of different circuit families implementing the same family of Boolean functions. In addition we analyze how heterogeneity in the underlying physical processes implementing the gates in a circuit influences the minimal thermodynamic cost of the overall circuit. These and other results of ours lay the foundation for extending circuit complexity theory to include mismatch cost as a resource cost.
@@ I Introduction
Fortunately, recent advances in non-equilibrium thermodynamics, particularly within the framework of stochastic thermodynamics, have made it possible to rigorously define and analyze thermodynamic quantities such as work, entropy, and heat dissipation in systems driven far from equilibrium van2015ensemble ; seifert2012stochastic . These tools are now being applied to fundamental questions about the thermodynamic costs of computation and communication processes van2015ensemble ; seifert2012stochastic ; wolpert2019stochastic ; wolpert2024thermodynamics ; wolpert2024stochastic . In particular, stochastic thermodynamics allows to analyze the entropy production (EP) of running any (physical system that implements) a computation, i.e., the unavoidably wasted energy incurred by running that computation.
@@
A central concept in this analysis is the mismatch cost (MMC) kolchinsky2021dependence ; kolchinsky2017dependence , which provides a lower bound on entropy production in physical computational processes.
@@
The structure of the paper is as follows: In Sec. II.1, we provide a minimal background on circuit complexity theory, reviewing the concepts of circuit families, size and depth complexity, and their associated complexity classes. In Sec. II.2, we briefly review stochastic thermodynamics and the mismatch cost lower bound on EP. In Sec. III, we focus on computing the mismatch cost for circuits. The circuit diagram alone does not specify the dynamical changes of the circuit’s state, i.e., of the joint distribution over the circuits gates as it runs. In Sec. III.1 we explain what other information is needed, and how to use it to calculate that change in the state of a circuit evolves during its execution. In Sec. III.2 and Sec. III.3 we combine these results, to derive the expression for the total MMC in computation with circuits.
Sec. IV builds on this earlier analysis to introduce mismatch cost complexity as a new circuit complexity measure, analogous to size and depth complexity, representing the minimal energetic cost of implementing a Boolean function with any circuit. We derive results and theorems that link MMC complexity to size complexity, showing that under certain conditions, the upper bound on MMC grows linearly with the size complexity of a circuit family. Complementing this result, we identify conditions where MMC scales differently from size complexity. In Sec.V, we discuss the energetic aspects of circuit optimization, comparing circuit families that implement a given Boolean function. Specifically, we study two families that compute the ADD function—one with depth complexity $\log(n)$ and another with $n$—to compare their mismatch costs. We also investigate the impact of heterogeneity in the physical properties of gates on the associated energetic costs. Finally, Sec. VI concludes the paper by discussing possible directions for future research.
@@ II.2 Stochastic Thermodynamics and Mismatch Cost
One of the central quantities in stochastic thermodynamics is entropy production (EP), which quantifies the irreversible heat dissipation in a process. If the probability distribution over the states of a physical system during its evolution changes from $p_{0}$ to $p_{\tau}$, the total EP associated with an initial distribution $p_{0}$ can be expressed as van2015ensemble ; kolchinsky2017dependence ,
@@ III.1 Dynamics of the circuit
$\mathcal{MC}_{\mu}$ Mismatch cost of running gate $\mu$ in a circuit
$\mathcal{MC}_{l}$ Mismatch cost of running layer $l$ in a circuit
$\mathcal{MC}_{ow}(p_{in})$ Mismatch cost of overwriting inputs with input distribution $p_{in}$
$\mathcal{MC}(C_{n},p_{in})$ Total mismatch cost of running the circuit $C_{n}$ for the input distribution $p_{in}$
@@ III.1.2 Joint distribution after the re-initialization of the input nodes
$p^{1}(\boldsymbol{x})=p_{in}(\boldsymbol{x}_{in})p(\boldsymbol{x}_{nin}).$ (10)
@@ III.1.3 Joint distribution after the update of gate $\mu$
III.2 Mismatch cost of re-initializing inputs with new values.
@@
The mismatch cost of overwriting the input values is
@@
Moreover, if $\mathrm{X}_{1}$ is a deterministic function of $\mathrm{X}_{in}$, then $\mathrm{I}_{0}(\mathrm{X}_{in};\mathrm{X}_{1})=S(p_{in})$. In that case, the overwriting mismatch cost is given by
@@
III.3 Mismatch cost of gate-by-gate or layer-by-layer implementation of the circuit.
@@
The mismatch cost of updating the gate $\mu$ is given by,
@@
Using equations (11), (12), (21), and (22), the expression for the mismatch cost $\mathcal{MC}_{\mu}$ simplifies to
@@
where $\mathrm{I}_{\mu}(\mathrm{X}_{\mu};\mathrm{X}_{\mathrm{ch}(\mu)})$ denotes the mutual information between gate $\mu$ and its children $\mathrm{ch}(\mu)$ prior to the update of $\mu$. Specifically, the mutual information is evaluated with respect to the distribution $p^{\mu}(\boldsymbol{x})$, which characterizes the state of the system immediately before the update of $\mu$. The derivation of Eq. (24) is provided in App. A.3. The total mismatch cost of executing the entire circuit through a sequential gate-by-gate implementation is given by the sum of the mismatch costs of individual gates and the mismatch costs incurred when updating input nodes,
@@
for any $l\in\{1,...,L\}$. Let $X_{l}$ be the random variable denoting the joint state of gates in layer $l$ and let $\boldsymbol{x}_{l}$ denote a value of $X_{l}$. In a layer-by-layer implementation, all the gates in a layer are updated simultaneously. Then, analogous to Eq. 24, the mismatch cost of updating layer $l$ in the circuit is given by,
@@
Updating multiple gates simultaneously in a layer-by-layer approach, instead of updating gates sequentially one by one, represents a form of time-coarse graining. This is because grouping several updates into a single time step effectively reduces the temporal resolution of the system’s dynamics. As demonstrated in yadav2024mismatch , mismatch cost decreases under time-coarse graining. Applied to circuits, this result implies that the mismatch cost of a layer-by-layer implementation always serves as a lower bound on the mismatch cost of a gate-by-gate implementation. In In Fig. 5, we compare the mismatch costs of the layer-by-layer and gate-by-gate implementations of the same ripple-carry adder. The gate-by-gate mismatch cost is computed using Eq. 24, while the layer-by-layer mismatch cost is computed using Eq. 29. As a consequence of the lowering in mismatch cost under time-coarse graining, the layer-by-layer mismatch cost is lower than the gate-by-gate mismatch cost for the same circuit, across ripple-carry adder circuits with varying input lengths.
Additionally, note that as the computation in a circuit progress—whether gate-by-gate or layer-by-layer—the distribution over the states cycles through a sequence of distributions (see Fig. 3 and 4). It is important to emphasize that, for the calculation of the total mismatch cost, when calculating the total mismatch cost, the phase at which the sequence of distributions is started does not impact the final mismatch cost, as long as the full cycle of execution is completed.
In the next section, we use Eq 24 to calculate the mismatch cost of various circuit families and broadly investigate how does the circuit structure affect their mismatch cost. In particular, we define mismatch cost complexity, analogous to the size and depth complexity of a circuit family. We will then establish results that relate mismatch cost complexity to size complexity.
Figure 5: Comparison of layer-by-layer and gate-by-gate mismatch costs for the ripple-carry adder circuit family. As the circuit size increases with the input size, the mismatch cost grows for both the layer-by-layer (dashed red) and gate-by-gate (solid blue) implementations. Notably, the layer-by-layer mismatch cost remains consistently lower than the gate-by-gate mismatch cost. In this comparison, the input distribution is uniformly random for both implementations, and the priors of all gates in the circuits are also uniform.
@@ IV The mismatch cost complexity of a circuit family
We aim to explore the relationship between the total mismatch cost of a circuit and other complexity measures, such as size and depth. To this end, we extend the definition of a basis for a circuit family to include the prior distribution associated with each logic gate in the basis. For a given basis $\mathcal{B}$, we define the associated prior basis ~=$\tilde{\mathcal{B}}=\{(g,q_{g,\mathrm{pa}(g)}):g\in\mathcal{B}\}$, where $q_{g,\mathrm{pa}(g)}$ denotes the prior distribution associated with gate $g\in\mathcal{B}$. We sometimes use $q_{g}$ as a shorthand for $q_{g,\mathrm{pa}(g)}$. The definition of extended prior basis allows us to define mismatch cost complexity associated with a circuit family.
@@
The mismatch cost complexity associated with any circuit family is influenced by several key properties of the circuit. First, it depends on the basis $\tilde{\mathcal{B}}$, which includes the prior associated with each allowed logic gate in the circuit. Second, the topology of the circuit, in combination with the input distribution, jointly determines the mutual information $\mathrm{I}_{\mu}(\mathrm{X}_{\mu};\mathrm{X}_{\mathrm{ch}(\mu)})$, which contributes to the mismatch cost for each gate $\mu$ downstream (see Eq. 24). Additionally, the topology and input distribution together determine the joint distribution $p_{\mu}(\boldsymbol{x}_{\mu})$ and $p_{\mu}(\boldsymbol{x}_{\mathrm{pa}}(\mu))$ for each gate $\mu$ downstream.
@@
The mismatch cost of running a gate $\mu$ in a circuit is upper bounded by
@@
Moreover, if $q_{\mu}$ is identical for all gates in all circuits in the circuit family, and $p_{in}^{n}$ is the input distribution for the circuit with input size $n$, then the total mismatch cost of any circuit $C_{n}$ in the circuit family is upper bounded by
@@
Thm. 1 establishes a relationship between the size complexity and the mismatch cost complexity of any circuit family constructed from gates in a homogeneous prior basis. If $s(n)$ denotes the size complexity of the circuit family, then according to Eq. (31), the mismatch cost complexity $m(n)$ of the circuit family is given by:
@@
Figure 7: Saturating the Upper Bound on MMC for the AND Binary Tree Circuit Family (Eq. (33)). When the input is always either all ones or all zeros, then $S(p^{n}_{in})=0$ and the theoretical upper bound is $n\ln\left(1\backslash q^{\min}\right)$. The solid blue line represents the total mismatch cost obtained by using Eq. 26 for an input distribution consisting entirely of either all ones or all zeros, matching exactly with the orange dashed line, which corresponds to the theoretical upper bound given by Eq. (33). In contrast, when the input distribution is a uniform distribution, the actual total mismatch cost of AND binary tree remains consistently lower than the theoretical upper bound: the solid green line depicts the total mismatch cost for a uniform input distribution, while the dashed red line shows its respective theoretical upper bound.
@@
As shown in Fig. 7, when the input is always either all ones or all zeros, the theoretical upper bound is fully attained by the actual total MMC of the circuit. However, for other input distributions, such as a uniform input distribution, the upper bound is not saturated, and the actual total mismatch cost remains consistently lower than the theoretical upper bound.
According to Eq. 32, if the size complexity $s(n)$ of a circuit family grows as a polynomial function of the input size, then the upper bound on the mismatch cost complexity for any homogeneous prior basis also scales polynomially with the input size. In other words, any Boolean function that can be implemented by a polynomial-size circuit can also be realized by a circuit family whose mismatch cost complexity does not exceed polynomial growth. This observation leads to the following corollary:
@@
However, it remains an open question whether the converse holds—that is, whether any function implementable by a circuit family with a mismatch cost that grows no faster than polynomial can also be realized by a circuit family whose size complexity is polynomially bounded, and thus whether $\mathbf{P_{/poly}}=\mathbf{MMC}(\mathrm{poly})$.
Next, we examine the scenario where the prior distribution is not homogeneous across the logic gates in the basis. For a given basis $\mathcal{B}$, we denote the prior distribution associated with each gate $g\in\mathcal{B}$ by $q_{g}$. The following result tells you how the mismatch cost complexity various with input size for an inhomogeneous prior basis:
@@
For a given basis $\mathcal{B}$, let $\{C_{n}\}_{n\in N}$ be circuit family and let and let $\tilde{\mathcal{B}}$ the associated prior basis. Then, the mismatch cost of circuit of input size $n$ is upper bounded by,
@@
The proof is provided in App. B.2. Note that Thm. 2 provides a slight generalization of Thm. 1 and applies to heterogeneous prior bases. Several key aspects distinguish it from the homogeneous prior case. For instance, in the homogeneous prior setting, the upper bound on the mismatch cost (MMC) is always proportional to the size complexity $|C_{n}|$ of the circuit family for every $n$. However, with heterogeneous priors, this proportionality may no longer hold.
To illustrate this, consider a circuit family $\{C_{n}\}_{n\in\mathbb{N}}$ whose basis comprises two types of gates — say, AND and XOR — labeled as $1$ and $2$. In this example, the number of AND gates scales as #$\#_{1}(n)=\log(n)$, while the number of XOR gates scales as #$\#_{2}(n)=n^{3}$. According to Thm. 2, the upper bound on the mismatch cost is given by:
@@
Now, suppose the priors associated with these two types of gates are such that $K_{1}\gg K_{2}$. Under this condition, for small $n$, the upper bound on the mismatch cost scales as $\log(n)$. However, as $n$ grows large, $n^{3}$ becomes the dominant term. This example demonstrates how, in the case of heterogeneous priors, the MMC can scale differently from the size complexity of the circuit family.
Figure 8: Mismatch cost scaling for a circuit family with homogeneous vs. heterogeneous priors. The circuit family that we consider here consist of binary tree made up AND and OR gates; each layer had one OR gate and rest of the gates are AND. The solid blue curve shows the mismatch cost when the prior of OR gates differs from the uniform prior of AND gates (heterogeneous priors). The mismatch cost is not linear with circuit size (Thm. 2). In contrast, the dashed orange curve represents the mismatch cost for the same circuit family with identical priors for XOR and AND gates (homogeneous priors), and the mismatch cost scales linearly with the circuit size.
@@
The lower bound on the MMC
Going back to Eq. (24) for the mismatch cost of a gate in a circuit, under the assumption that the prior distribution $q_{\mu,\mathrm{pa}(\mu)}(\boldsymbol{x}_{\mu},\mathrm{pa}(\mu))$ is a product distribution, $q_{\mu}(\boldsymbol{x}_{\mu})q_{\mathrm{pa}(\mu)}(\boldsymbol{x}_{\mathrm{pa}(\mu)})$, the expression for mismatch cost simplifies to:
@@
It turns out that for certain kinds of circuits, the mutual information between the gate and its children is zero, e.g., in an XOR binary tree. Moreover, if the actual distribution $p_{\mu}$ perfectly matches the prior $q_{\mu}$, the KL-divergence term becomes zero. Therefore, it is theoretically possible for the mismatch cost of a gate to approach zero.
This bound can even be achieved for certain types of circuits with specific prior distributions. For example, in an XOR binary tree fed with a uniform input distribution, the distribution $p_{\mu}$ remains a uniform distribution for every downstream XOR gate in the circuit. If all the XOR gates have a uniform prior, $q_{\mu}$, then the actual distribution $p_{\mu}$ perfectly matches with the prior, leading to zero KL divergence and, hence, zero mismatch cost.
However, even if the mismatch cost of each gate in the circuit is zero, the total mismatch cost includes the overwriting mismatch cost, which is lower bounded by the mutual information between input and non-input gates. Therefore, the total associated mismatch cost is lower bounded by the mutual information between the input and non-input gates,

### v2

@@ Abstract
All digital devices have components that implement Boolean functions, mapping that component’s input to its output. However, any fixed Boolean function can be implemented by an infinite number of circuits, all of which vary in their resource costs. This has given rise to the field of circuit complexity theory, which studies the minimal resource cost to implement a given Boolean function with any circuit. Traditionally, circuit complexity theory has focused on the resource costs of a circuit’s size (its number of gates) and depth (the longest path length from the circuit’s input to its output). In this paper, we extend circuit complexity theory by investigating the minimal thermodynamic cost of a circuit’s operation. We do this by using the mismatch cost (MMC) of a given circuit that is run multiple times in a row to calculate a lower bound on the entropy production (EP) incurred in each such run. Specifically, we discuss conditions under which MMC of a circuit is proportional to the size of a circuit, and conditions for them to diverge. We also use our results to compare the thermodynamic costs of different circuit families implementing the same family of Boolean functions. In addition we analyze how heterogeneity in the underlying physical processes implementing the gates in a circuit influences the minimal thermodynamic cost of the overall circuit. These and other results of ours lay the foundation for extending circuit complexity theory to include mismatch cost as a resource cost.
@@ I Introduction
Fortunately, recent advances in non-equilibrium thermodynamics, particularly within the framework of stochastic thermodynamics, have made it possible to rigorously define and analyze thermodynamic quantities such as work, entropy, and heat dissipation in systems driven far from equilibrium van2015ensemble ; seifert2012stochastic . These tools are now being applied to fundamental questions about the thermodynamic costs of computation and communication processes van2015ensemble ; seifert2012stochastic ; wolpert2019stochastic ; wolpert2024thermodynamics ; wolpert2024stochastic . In particular, stochastic thermodynamics allows us to analyze the entropy production in a process, i.e., the unavoidably wasted energy incurred by running that process.
@@
A central concept in this analysis is the mismatch cost (MMC) kolchinsky2021dependence ; kolchinsky2017dependence , which is a component of the entropy production arising in physical computational processes.
@@
The structure of the paper is as follows: In Sec. II.1, we provide a minimal background on circuit complexity theory, reviewing the concepts of circuit families, size and depth complexity, and their associated complexity classes. In Sec. II.2, we briefly review stochastic thermodynamics and the MMC lower bound on EP. In Sec. III, we focus on determining MMC of a circuit. The circuit diagram alone does not specify the dynamical changes of the circuit’s state, i.e., of the joint distribution over the circuits gates as it runs. In Sec. III.1 we explain what other information is needed, and how to use it to calculate that change in the state of a circuit evolves during its execution. In Sec. III.2 and Sec. III.3 we combine these results, to derive the expression for the total MMC in computation with circuits.
Sec. IV builds on this earlier analysis to introduce mismatch cost complexity as a new circuit complexity measure, analogous to size and depth complexity, representing the minimal energetic cost of implementing a Boolean function with any circuit. We derive results and theorems that link MMC complexity to size complexity, showing that under certain conditions, the upper bound on MMC grows linearly with the size complexity of a circuit family. Complementing this result, we identify conditions where MMC scales differently from size complexity. In Sec.V, we discuss the energetic aspects of circuit optimization by comparing the MMC associated with circuit families that implement a given Boolean function. Specifically, we study two families that compute the ADD function—one with depth complexity $\log(n)$ and another with $n$. We also investigate the impact of heterogeneity in the physical properties of gates on the associated MMC. Finally, Sec. VI concludes the paper by discussing possible directions for future research.
@@ II.2 Stochastic Thermodynamics and Mismatch Cost
One of the central quantities in stochastic thermodynamics is entropy production, which quantifies the irreversible heat dissipation in a process. If the probability distribution over the states of a physical system during its evolution changes from $p_{0}$ to $p_{\tau}$, the total EP associated with an initial distribution $p_{0}$ can be expressed as van2015ensemble ; kolchinsky2017dependence ,
@@ III.1 Dynamics of the circuit
$\mathcal{MC}_{\mu}$ MMC of running gate $\mu$ in a circuit
$\mathcal{MC}_{l}$ MMC of running layer $l$ in a circuit
$\mathcal{MC}_{ow}(p_{in})$ MMC of overwriting inputs with input distribution $p_{in}$
$\mathcal{MC}(C_{n},p_{in})$ Total MMC of running the circuit $C_{n}$ for the input distribution $p_{in}$
@@ III.1.2 Joint distribution after the re-initialization of the input nodes
$p^{1}(\boldsymbol{x})=p_{in}(\boldsymbol{x}_{in})p_{nin}(\boldsymbol{x}_{nin}).$ (10)
@@ III.1.3 Joint distribution after the update of gate $\mu$
III.2 MMC of re-initializing inputs with new values.
@@
The MMC of overwriting the input values is
@@
Moreover, if $\mathrm{X}_{1}$ is a deterministic function of $\mathrm{X}_{in}$, then $\mathrm{I}_{0}(\mathrm{X}_{in};\mathrm{X}_{1})=S(p_{in})$. In that case, the overwriting MMC is given by
@@
III.3 MMC of gate-by-gate or layer-by-layer implementation of the circuit.
@@
The MMC of updating the gate $\mu$ is given by,
@@
Using equations (11), (12), (21), and (22), the expression for the MMC $\mathcal{MC}_{\mu}$ simplifies to
@@
where $\mathrm{I}_{\mu}(\mathrm{X}_{\mu};\mathrm{X}_{\mathrm{ch}(\mu)})$ denotes the mutual information between gate $\mu$ and its children $\mathrm{ch}(\mu)$ prior to the update of $\mu$. Specifically, the mutual information is evaluated with respect to the distribution $p^{\mu}(\boldsymbol{x})$, which characterizes the state of the system immediately before the update of $\mu$. The derivation of Eq. (24) is provided in App. A.3. The total MMC of executing the entire circuit through a sequential gate-by-gate implementation is given by the sum of the MMCs of individual gates and the MMCs incurred when updating input nodes,
@@
for any $l\in\{1,...,L\}$. Let $X_{l}$ be the random variable denoting the joint state of gates in layer $l$ and let $\boldsymbol{x}_{l}$ denote a value of $X_{l}$. In a layer-by-layer implementation, all the gates in a layer are updated simultaneously. Then, analogous to Eq. 24, the MMC of updating layer $l$ in the circuit is given by,
@@
Updating multiple gates simultaneously in a layer-by-layer approach, instead of updating gates sequentially one by one, represents a form of time-coarse graining. This is because grouping several updates into a single time step effectively reduces the temporal resolution of the system’s dynamics. As demonstrated in yadav2024mismatch , MMC decreases under time-coarse graining. Applied to circuits, this result implies that the MMC of a layer-by-layer implementation always serves as a lower bound on the MMC of a gate-by-gate implementation. In Fig. 5, we compare the MMCs of the layer-by-layer and gate-by-gate implementations of the same ripple-carry adder. The gate-by-gate MMC is computed using Eq. 24, while the layer-by-layer MMC is computed using Eq. 29. As a consequence of the lowering in MMC under time-coarse graining, the layer-by-layer MMC is lower than the gate-by-gate MMC for the same circuit, across ripple-carry adder circuits with varying input lengths.
Additionally, note that as the computation in a circuit progress—whether gate-by-gate or layer-by-layer—the distribution over the states cycles through a sequence of distributions (see Fig. 3 and 4). It is important to emphasize that, for the calculation of the total MMC, when calculating the total MMC, the phase at which the sequence of distributions is started does not impact the final MMC, as long as the full cycle of execution is completed.
Moreover, the sequence of gate implementations in a circuit can be modeled as a periodic process. In practical computational devices based on circuits, the execution of gates or layers is typically driven by a clock. From this perspective, the process is periodic, with gates executed at each clock cycle, sequentially leading to the completion of the full circuit. In Appendix B.3, we present the MMC calculation for this clock-driven model of circuit execution. This approach yields the same expression for the MMC as given in Eq. 23.
In the next section, we use Eq 24 to calculate the MMC of various circuit families and broadly investigate how does the circuit structure affect their MMC. In particular, we define MMC complexity, analogous to the size and depth complexity of a circuit family. We will then establish results that relate MMC complexity to size complexity.
Figure 5: Comparison of layer-by-layer and gate-by-gate mismatch costs for the ripple-carry adder circuit family. As the circuit size increases with the input size, the MMC grows for both the layer-by-layer (dashed red) and gate-by-gate (solid blue) implementations. Notably, the layer-by-layer MMC remains consistently lower than the gate-by-gate MMC. In this comparison, the input distribution is uniformly random for both implementations, and the priors of all gates in the circuits are also uniform.
@@ IV The mismatch cost complexity of a circuit family
We aim to explore the relationship between the total MMC of a circuit and other complexity measures, such as size and depth. To this end, we extend the definition of a basis for a circuit family to include the prior distribution associated with each logic gate in the basis. For a given basis $\mathcal{B}$, we define the associated prior basis ~=$\tilde{\mathcal{B}}=\{(g,q_{g,\mathrm{pa}(g)}):g\in\mathcal{B}\}$, where $q_{g,\mathrm{pa}(g)}$ denotes the prior distribution associated with gate $g\in\mathcal{B}$. We sometimes use $q_{g}$ as a shorthand for $q_{g,\mathrm{pa}(g)}$. The definition of extended prior basis allows us to define MMC complexity associated with a circuit family.
@@
The MMC complexity associated with any circuit family is influenced by several key properties of the circuit. First, it depends on the basis $\tilde{\mathcal{B}}$, which includes the prior associated with each allowed logic gate in the circuit. Second, the topology of the circuit, in combination with the input distribution, jointly determines the mutual information $\mathrm{I}_{\mu}(\mathrm{X}_{\mu};\mathrm{X}_{\mathrm{ch}(\mu)})$, which contributes to the MMC for each gate $\mu$ downstream (see Eq. 24). Additionally, the topology and input distribution together determine the joint distribution $p_{\mu}(\boldsymbol{x}_{\mu})$ and $p_{\mu}(\boldsymbol{x}_{\mathrm{pa}}(\mu))$ for each gate $\mu$ downstream.
@@
The MMC of running a gate $\mu$ in a circuit is upper bounded by
@@
Moreover, if $q_{\mu}$ is identical for all gates in all circuits in the circuit family, and $p_{in}^{n}$ is the input distribution for the circuit with input size $n$, then the total MMC of any circuit $C_{n}$ in the circuit family is upper bounded by
@@
Thm. 1 establishes a relationship between the size complexity and the MMC complexity of any circuit family constructed from gates in a homogeneous prior basis. If $s(n)$ denotes the size complexity of the circuit family, then according to Eq. (31), the MMC complexity $m(n)$ of the circuit family is given by:
@@
Figure 7: Saturating the Upper Bound on MMC for the AND Binary Tree Circuit Family (Eq. (33)). When the input is always either all ones or all zeros, then $S(p^{n}_{in})=0$ and the theoretical upper bound is $n\ln\left(1\backslash q^{\min}\right)$. The solid blue line represents the total MMC obtained by using Eq. 26 for an input distribution consisting entirely of either all ones or all zeros, matching exactly with the orange dashed line, which corresponds to the theoretical upper bound given by Eq. (33). In contrast, when the input distribution is a uniform distribution, the actual total MMC of AND binary tree remains consistently lower than the theoretical upper bound: the solid green line depicts the total MMC for a uniform input distribution, while the dashed red line shows its respective theoretical upper bound.
@@
As shown in Fig. 7, when the input is always either all ones or all zeros, the theoretical upper bound is fully attained by the actual total MMC of the circuit. However, for other input distributions, such as a uniform input distribution, the upper bound is not saturated, and the actual total MMC remains consistently lower than the theoretical upper bound.
According to Eq. 32, if the size complexity $s(n)$ of a circuit family grows as a polynomial function of the input size, then the upper bound on the MMC complexity for any homogeneous prior basis also scales polynomially with the input size. In other words, any Boolean function that can be implemented by a polynomial-size circuit can also be realized by a circuit family whose MMC complexity does not exceed polynomial growth. This observation leads to the following corollary:
@@
However, it remains an open question whether the converse holds—that is, whether any function implementable by a circuit family with a MMC that grows no faster than polynomial can also be realized by a circuit family whose size complexity is polynomially bounded, and thus whether $\mathbf{P_{/poly}}=\mathbf{MMC}(\mathrm{poly})$.
Next, we examine the scenario where the prior distribution is not homogeneous across the logic gates in the basis. For a given basis $\mathcal{B}$, we denote the prior distribution associated with each gate $g\in\mathcal{B}$ by $q_{g}$. The following result tells you how the MMC complexity various with input size for an inhomogeneous prior basis:
@@
For a given basis $\mathcal{B}$, let $\{C_{n}\}_{n\in N}$ be circuit family and let and let $\tilde{\mathcal{B}}$ the associated prior basis. Then, the MMC of circuit of input size $n$ is upper bounded by,
@@
The proof is provided in App. B.2. Note that Thm. 2 provides a slight generalization of Thm. 1 and applies to heterogeneous prior bases. Several key aspects distinguish it from the homogeneous prior case. For instance, in the homogeneous prior setting, the upper bound on the MMC is always proportional to the size complexity $|C_{n}|$ of the circuit family for every $n$. However, with heterogeneous priors, this proportionality may no longer hold.
To illustrate this, consider a circuit family $\{C_{n}\}_{n\in\mathbb{N}}$ whose basis comprises two types of gates — say, AND and XOR — labeled as $1$ and $2$. In this example, the number of AND gates scales as #$\#_{1}(n)=\log(n)$, while the number of XOR gates scales as #$\#_{2}(n)=n^{3}$. According to Thm. 2, the upper bound on the MMC is given by:
@@
Now, suppose the priors associated with these two types of gates are such that $K_{1}\gg K_{2}$. Under this condition, for small $n$, the upper bound on MMC scales as $\log(n)$. However, as $n$ grows large, $n^{3}$ becomes the dominant term. This example demonstrates how, in the case of heterogeneous priors, the MMC can scale differently from the size complexity of the circuit family.
Figure 8: Mismatch cost scaling for a circuit family with homogeneous vs. heterogeneous priors. The circuit family that we consider here consist of binary tree made up AND and OR gates; each layer had one OR gate and rest of the gates are AND. The solid blue curve represents MMC when the prior of OR gates differs from the uniform prior of AND gates (heterogeneous priors). MMC is not linear with circuit size (Thm. 2). In contrast, the dashed orange curve represents the MMC for the same circuit family with identical priors for XOR and AND gates (homogeneous priors), and MMC scales linearly with the circuit size.
Figure 9: Mismatch cost for binary trees whose internal nodes are all 2-input AND gates, shown for depths 1–4. The three curves correspond to distinct designer priors on the gate’s input–output triple $(x,y,g)$. The agreeing prior places 97 % probability on the output that an ideal AND would produce; the uniform prior assigns equal mass to all eight states; the disagreeing prior places 97 % probability on the output opposite to AND. This clearly shows the slope of the cost vs size line depends on the prior
Figure 10: This figure illustrates that the linear cost–size relation need not hold for circuit families in the finite input regime. The figure is made by creating circuit families by “mixing and matching” small perfect-tree circuits. The first point is a depth-1 tree. For the depth-2 tree (three internal gates) we pick the truth table with the smallest per-gate slope under a common skewed prior $Q_{par}(0,0)=0.9$,$Q_{g}(0)=0.9$. The depth-3 tree uses the second-smallest-slope gate, and the depth-4 tree uses the gate with the largest slope. Although each individual segment is linear in circuit size, the deliberate change of gate type at each depth yields a piecewise-linear curve whose overall shape is distinctly non-linear, illustrating how mixing gate families can steer the global cost profile even when the prior is held fixed for all gates
@@
The lower bound on the mismatch cost
Going back to Eq. (24) for MMC of a gate in a circuit, under the assumption that the prior distribution $q_{\mu,\mathrm{pa}(\mu)}(\boldsymbol{x}_{\mu},\mathrm{pa}(\mu))$ is a product distribution, $q_{\mu}(\boldsymbol{x}_{\mu})q_{\mathrm{pa}(\mu)}(\boldsymbol{x}_{\mathrm{pa}(\mu)})$, the expression for MMC simplifies to:
@@
It turns out that for certain kinds of circuits, the mutual information between the gate and its children is zero, e.g., in an XOR binary tree. Moreover, if the actual distribution $p_{\mu}$ perfectly matches the prior $q_{\mu}$, the KL-divergence term becomes zero. Therefore, it is theoretically possible for MMC of a gate to approach zero.
This bound can even be achieved for certain types of circuits with specific prior distributions. For example, in an XOR binary tree fed with a uniform input distribution, the distribution $p_{\mu}$ remains a uniform distribution for every downstream XOR gate in the circuit. If all the XOR gates have a uniform prior, $q_{\mu}$, then the actual distribution $p_{\mu}$ perfectly matches with the prior, leading to zero KL divergence and, hence, zero MMC.
However, even if the MMC of each gate in the circuit is zero, the total MMC includes the overwriting mismatch cost, which is lower bounded by the mutual information between input and non-input gates. Therefore, total MMC is lower bounded by the mutual information between the input and non-input gates,